OpenStudy (anonymous):

how to do 7x - 3y = 20 ---------- 5x + 3y = 16 with elimination?

OpenStudy (anonymous):

the 3y cancels out, so 7x+5x=12x and 20+16= 36so the new equation is 12x=36

OpenStudy (anonymous):

add two equations and u will get 12x=36 so x=3, put this value of x in any equation, u will get the value of y

OpenStudy (anonymous):

I thought with elimination you were supposed to multiply one of the equations to make it the same as the other? Im not completely sure thats why i asked here :P

OpenStudy (anonymous):

Well, you do need to multiply one or both equations to make the coefficient of one of the variables same (constant before the variable). In this case, the coefficient of y is already the same (3). So, you don't need to multiply to do elimination.

OpenStudy (anonymous):

If you had for example x + 2y = 5 2x + y = 7 Then, you will have to multiply the second equation by 2 to eliminate y. Or, multiply the first equation by 2 to eliminate x.

OpenStudy (anonymous):

good work